# Dom Huh    Prasant Mohapatra

### v1

@@ Abstract
This paper proposes a diffusion-based auto-bidding framework that leverages graph representations to model large-scale auction environments. In such settings, agents must dynamically optimize bidding strategies under constraints defined by key performance indicator (KPI) metrics, all while operating in competitive environments characterized by uncertain, sparse, and stochastic variables. To address these challenges, we introduce a novel approach combining learnable graph-based embeddings with a planning-based latent diffusion model (LDM). By capturing patterns and nuances underlying the interdependence of impression opportunities and the multi-agent dynamics of the auction environment, the graph representation enable expressive computations regarding auto-bidding outcomes. With reward alignment techniques, the LDM’s posterior is fine-tuned to generate auto-bidding trajectories that maximize KPI metrics while satisfying constraint thresholds. Empirical evaluations on both real-world and synthetic auction environments demonstrate significant improvements in auto-bidding performance across multiple common KPI metrics, as well as accuracy in forecasting auction outcomes.
®ate Optimization
1 Introduction
Optimal auto-bidding in large-scale auctions remains an open research challenge with significant real-world implications for online markets. In these environments, auto-bidding agents must compete with bids on high-volume, stochastic streams of impression opportunities (IOs) to maximize value while adhering to multiple key performance indicator (KPI) metric constraints (Aggarwal et al., 2024; Kitts et al., 2017). Recent efforts (Guo et al., 2024) have demonstrated impressive successes with using generative diffusion models (Sohl-Dickstein et al., 2015) to simulate auction trajectories and plan bids to maximize their KPI adherence. However, existing methods rely on heuristic feature representations that inadequately capture the interdependencies between IOs and the multi-agent dynamics of auction environments.
@@ 1 Introduction
Figure 1: An illustration of the LGD-AB framework. At every time-step, a graph is formed for agent $i$. The nodes of a sub-graph $\mathcal{G}_{t}^{i}$ represent the IO of agent $i$. Sub-graphs for other agents are similarly formed and connected accordingly. The graph is processed using a GNN to generate an embedding vector $x^{i}_{t}$. The diffusion model then learns a posterior of the sequence of embedding vectors.
To address this discrepancy, we introduce a graph-based representation of auto-bidding dynamics, modeling IOs on a more granular level to retain their individual attributes as well as encoding their connections to other IOs and agents. We then extend latent diffusion models (LDMs) (Rombach et al., 2022) to leverage this structure, formulating our framework – denoted Latent Graph Diffusion Model for Auto-Bidding (LGD-AB) – to forecast auction outcomes and extract crucial bidding information from a more expressive feature space. Further, we align the LDM (Huh & Mohapatra, 2025) to optimize against the KPI metrics. Our empirical experiments on real-world and synthetic auction simulations (Su et al., 2025; Deng et al., 2024) demonstrate that LGD-AB outperforms existing methods across common KPIs (He et al., 2021).
2 Latent Graph Diffusion Model for Auto-Bidding
@@
2.1 Graph Embedding For Ad Auctions
The objective of the graph embedding is to represent the IOs presented to each agent in an expressive and comprehensive manner. In large-scale auctions, the number of IOs exposed to each agent can range drastically, from 00 to $10$s of thousands at every time-step. To process data with such variability yet interdependency, we represent these IOs as nodes of a graph and process their features using graph convolutions with an attention-based graph neural network (GNN) (Brody et al., 2022). For scalability and computational limitations for larger scale simulations, we cluster IOs to avoid prohibitively large graphs. We additionally allocate a node dedicated for non-exposed IOs to maintain context during sparse time periods. In this graph representation, we fully connect the nodes belonging to same agents.
We optimize the latent graph embedding $x^{i}_{t}$ of agent $i$ at time-step $t$ generated by the GNN to hold information regarding the bids $b_{i}^{t}[k]$ made for each IO $k$. Hence, we introduce a separate inverse dynamics model (IDM) to support the optimization of the following objective:
$\mathcal{L}_{\text{graph}}(\theta)=\mathbb{E}_{(t,k)}[\log\text{P}(b_{t}^{i}[k]|\theta,x_{t}^{i},f_{t}^{i}[k],c_{t}^{i})]$ (1)
where $k$ is the IO identification index and $f_{t}^{i}[k]$ are the node features of the corresponding IO. To ensure the feasibility of this optimization, we append a context vector $c_{t}^{i}$ that provides more information regarding the agent agnostic to the individual IO. To clarify, we optimize the objective over the parameters $\theta$ of both the GNN and the IDM. We view this optimization to (1) instill bid-policy information into the embedding and (2) constrain the graph embedding to reflect auction dynamics.
Graphs of Other Agents
To ensure the embedding space accounts for the presence of other agents within the auction setting, we define additional sub-graphs representing the other agents constructed identically. The nodes between agents that share IOs are connected to encode the relational information that the exposures of IO can be shared between agents. IO nodes that are not exposed to an agent are connected to the unique IO node dedicated to non-exposed IOs, as seen in Figure 1.
Figure 2: Agent $i$’s approximation of Other Agents’ Graph $\omega^{i}(\mathcal{G}_{t}^{-i})$, with two nodes for student GNN.
@@ 2.1 Graph Embedding For Ad Auctions
We note that the advertiser’s visibility of other agents’ IOs and bids is contingent to the auction design and prevents the full construction of the graphs of other agents. In such cases, we extend the concept of centralized training and decentralized execution paired with knowledge distillation (KD) (Gou et al., 2021). We train a centralized GNN $\mathcal{T}$ with full visibility of all agents’ sub-graphs on Equation 1. We then train a student model $\mathcal{S}$ using a response-based KD loss to mimic $\mathcal{T}$. As shown in Figure 2, $\mathcal{S}$ approximates the other agents’ sub-graph with two proxy nodes with learnable node feature embeddings. Similarly, if the non-exposed IOs are not visible to the agent, we learn its node feature embedding.
@@
We extend representation learning techniques, namely MA-SPL (Huh & Mohapatra, 2024), to further embed multi-agent and temporal dynamics information within the latent graph embedding $\{x^{i}_{0},x_{1}^{i},\dots,x_{T}^{i}\}_{i\in A}$, where $A$ is the set of all agents.
Approximate Equilibrium Computation
To promote a more cohesive multi-agent bidding strategy, we compute a joint embedding space by merging all agents’ graph embeddings and we refer to this design choice as approximate equilibrium computation (EC). Therefore, we introduce an aggregation operation $\mathcal{A}:\{x^{i}_{t}\}_{i\in A}\mapsto x_{t}$ using a self-attention network. This latent embedding $x_{t}$ is now shared between all agents when computing the embedding and generative objectives as well as for individual bid formulation.
2.2 Latent Diffusion Models
The latent diffusion model (LDM) processes a temporal sequence of graph embeddings $\{x_{t}\}_{t\in[\tau,\tau+T]}$, where $T$ is the length of the context window, and forms the foundation to devise a planning-based auto-bidding solution. In this work, we adopt the Decision Diffuser framework (Ajay et al., 2023), denoising state trajectories and offloading the action generation to the IDM. Unlike recent works using transformer-based architectures (Peebles & Xie, 2023) and score-based prediction (Song et al., 2020), we found greater success using the traditional 1D temporal convolution network with $\epsilon$-prediction. We optimize the diffusion objective using in-painting masks (Lugmayr et al., 2022).
@@ 2.2 Latent Diffusion Models
To ensure that the LDM’s posterior aligns with the auto-bidding objective, we follow the procedure set in (Huh & Mohapatra, 2025), utilizing an iterative sequence of constrained alignment optimization of reinforcement learning, direct preference optimization, and supervised fine-tuning with rejection sampling. We train the value function using IQL (Kostrikov et al., 2021) to predict not only the return, i.e., the number of total conversations, but also other important KPI metrics, such as CPA, ROI, CVR, budget adherence, and social welfare. We define the conversion rate, i.e., CVR, as the ratio between the number of winning bids and the number of total bids made. With this value function, we can strictly enforce KPI constraints, most notably, budget adherence, through rejection sampling.
@@
Training LGD-AB occurred in two stages. We first trained the graph embedding module and then the LDM separately. We then train both modules together, fine-tuning the GNN at a lowered learning rate. We found that this additional stage of joint training benefits the performance of both modules.
3 Results and Ablations Studies
3.1 Experiment Setup
We evaluate LGD-AB on two auto-bidding simulations: AuctionNet (Su et al., 2025) and synthetic auction (Deng et al., 2024).
AuctionNet (Su et al., 2025) contains over $500$ million in real-world records of online advertising instances, with $21$ separate bidding periods of $48$ time-steps and $48$ different advertisers. We constructed an auto-bidding simulator using this dataset, following prior works (Chiappa et al., 2024).
Synthetic auction, on the other hand, describes a sound procedure to generate artificial auction data that mimics real ad auctions by capturing its underlying properties through hierarchical feature generation and specific probabilistic distributions. We extend (Deng et al., 2024) with configurable parameters to introduce longer and variable ad life cycles, multi-slot winning IOs, and a budget constraint on each agent. We collect an offline RL dataset of $100$ thousand auction simulations with a varying number of advertisers, allocation rule (e.g., FPA, GSP, VCG), bidding period durations, and IO’s properties. Within this dataset, advertisers either follow a uniform bid-scaling strategy or the non-uniform bid-scaling algorithm, as described in (Deng et al., 2024).
3.2 Empirical Evaluation
Auction Forecasting
$F(\theta)=\mathbb{E}_{x,t}[\log\text{P}(x_{t^{\prime}>t}|\theta,x_{t^{\prime}\leq t})]$ (2)
In Table 1, our evaluation was performed using our online simulators, where to compute $F(\theta)$, we use a testing dataset that was not used during training. We found, with KD, the student model retains on average $77.48\%$ of the teacher model’s forecasting accuracy, despite partial observability. Interestingly, EC improves forecasting capability in Synthetic Auction, however, in AuctionNet, it failed to replicated this improvement. We attribute this to the limited number of data samples (i.e., only $21$ bidding periods in total), we utilize synthetic data in AuctionNet to mitigate the data scarity. After further training, we found a more fair representation of its forecasting accuracy. Lastly, the KPI alignment reduced average forecasting accuracy by $9.47\%$, which is expected, as such fine-tuning may lower the likelihood of lower value auto-bidding outcomes.
$F(\theta)$
AuctionNet Synthetic Auction
$\mathcal{T}$ $3.43$ $2.59$
$\mathcal{S}$ $2.64$ $2.02$
No EC $3.37$ $2.13$
EC $0.21\rightarrow 3.32^{*}$ $2.62$
No Alignment $3.25$ $2.09$
Alignment $3.02$ $2.04$
Table 1: Auction Forecasting Results.
@@ 3.2 Empirical Evaluation
Table 2: Mean and ±$\pm 1$ std performances between DiffBid and LGD-AB ($\mathcal{S}$, EC, and post-alignment) over $64$ random seeds.
Figure 3: KPI Alignment Learning Curve on Synthetic Auction with multi-KPI alignment compared to optimizing each KPI criterion independently. For social welfare, we compare the training between EC and non-EC variants.
@@
We benchmark our LGD-AB to a competitive diffusion-based planning solution, DiffBid (Guo et al., 2024) across six common KPI metrics. We highlight that the main differences of the DiffBid framework to our LGD-AB are its use of a heuristic feature space, bid-scaling, and a conditional return input. Comparisons to other generative models (i.e., Decision Transformer) and RL baselines (i.e., BC, CQL, and IQL) to DiffBid are provided in (Guo et al., 2024), where DiffBid has demonstrated notable improvements.
From Table 2, we observe improvements across all KPIs using our LGD-AB framework. The results demonstrate a significant increase in the number of IOs won, shown by average ×$\times 1.29$ improvement in return from our competitive baseline while significantly maximizing CVR and lowering CPA under adherence to the budget constraint. Lastly, our proposed LGD-AB has demonstrated better social welfare, indicating a collective utility to the auction. In both our baseline and LGD-AB solution, budget adherence was not an issue and remained comparable, maintaining near-perfect budget adherence over all trials.
As shown in Figure 3, we find that optimizing over multiple KPI (i.e., return, CPA, ROI, CVR, and social welfare) at the same time provides meaningful improvements within the individual performances of each metric. Additionally, the results affirm the utility of EC, as having a shared latent graph embedding can enable proper and more effective optimization of social welfare.
Accuracy of Bid Computation
We compare the bid prediction accuracy of DiffBid and LGD-AB in Table 3, depicting an improvement in being able to extract bids more accurately from our graph embeddings. Notably, in the Synthetic Auction simulation dataset, where portions of the auction data are generated from non-uniform bid-scaling strategies, we notice a wider margin of difference in accuracy in favor of our LGD-AB framework.
@@
Table 3: Comparsion between Bid Computation Accuracy of DiffBid and LGD-AB.
@@
We observe two key issues and limitations while developing our proposed solution, mainly pertaining to the scale of data and efficiency of our generative modeling algorithm. With AuctionNet, while there exists a large set of individualized auction records to train on, there was an insufficient number of collective auction records that span full bidding periods to properly train a generative model to capture the underlying auction dynamics accurately. We addressed this limitation with synthetic data generation and hyper-parameter tuning on the diffusion model. However, this limited performance can be improved with advancements in either training generative models in low data regimes (e.g., few-shot learning) or greater data availability. The other issue concerns the computational limitation from scaling the graph representation w.r.t. the number of IOs exposed to each agent. In theory, the original proposition was to allocate a node for each IO with edges denoting various aspects of their pairwise inter-relations, which scales poorly with a large number of shared IOs; an important feature was noticed in AuctionNet. Again, while we managed to address this with a clustering approach for AuctionNet, this reduced representation significantly reduced the expressiveness of the graph embeddings. In future works, a form of dynamic graph sparsification or hierarchical graph representations could mitigate this.
4 Conclusion
In this work, we propose a latent diffusion planning framework using graph-based embedding for auto-bidding applications. We demonstrate the utility of a graph-based representation, which enabled the LDM to capture the underlying dynamics distribution of a multi-agent auction process. Through aligning various KPIs on the trained LDM, we devised an effective bidding strategy capable of generating bids that maximize various KPI metrics.

### v3

# Davis, California, United States

@@ Abstract
This paper presents a novel diffusion-based auto-bidding framework that uses graph representations to model large-scale auction environments. In such environments, agents must optimize bidding strategies dynamically, balancing key performance indicators (KPIs) while navigating competitive, uncertain, and sparse conditions. To address these challenges, we introduce an approach that combines learnable graph embeddings with a planning-based latent diffusion model (LDM). This model captures the intricate relationships between impression opportunities and multi-agent interactions, enabling more accurate predictions of auto-bidding outcomes. Through reward-alignment techniques, the LDM’s posterior is fine-tuned to maximize KPI performance under predefined constraints. Our experiments, conducted in both real-world and synthetic auction environments, show significant improvements in the accuracy of auction outcome forecasts through learnable graph-based embeddings and in bidding performance across a range of common KPIs.
Index Terms:
auction, auto-bidding, graph, diffusion model
[…]
Preprint
I Introduction
Optimal auto-bidding in large-scale auctions remains an open research challenge with significant real-world implications for online markets, particularly in domains like digital advertising. In these dynamic auction environments, auto-bidding agents compete by placing bids on high-volume, stochastic streams of impression opportunities (IOs). The objective is to maximize the value of the bids while adhering to multiple key performance indicators (KPI) constraints, such as budget adherence, return on investment (ROI), conversion rates (CVR), social welfare, and cost-per-acquisition (CPA) [1, 2]. The inherent uncertainty and competition in these auctions make the optimization task particularly difficult [3], as agents must not only understand these interactions but also constantly adapt their strategies to account for changing conditions.
Recent advancements have shown the potential of generative diffusion models [4] to simulate auction trajectories and plan bids, with a focus on maximizing KPI adherence [5]. These methods have demonstrated success in capturing complex, high-dimensional patterns in bidding strategies. However, current approaches primarily rely on heuristic feature representations that fail to capture the full extent of the interdependencies between IOs and the dynamic, multi-agent nature of auction environments. These limitations often hinder the ability to model the intricate relationships between agents, auction parameters, and impression opportunities, leading to suboptimal bidding strategies.
@@ I Introduction
Figure 1: An illustration of the LGD-AB framework. At every time-step, a graph-based embedding is computed for agent $i$ using a bipartite graph of agent and IO nodes. The nodes of a sub-graph $\mathcal{G}_{t}^{i}$ consist of two virtual nodes, connecting to IOs exposed and not exposed to agent $i$ respectively. Sub-graphs for other agents are similarly formed and connected accordingly. The graph is processed using a GNN to generate an embedding vector $x^{i}_{t}$. The diffusion model then learns a posterior of the temporal sequence of embedding vectors to forecast auction dynamics.
To address these challenges, we propose a novel framework that incorporates graph-based representations to model auction environments and a latent diffusion model (LDM) to plan and optimize bidding strategies. Our framework, Latent Graph Diffusion Model for Auto-Bidding (LGD-AB), goes beyond traditional heuristics by using learnable graph embeddings to represent the complex interrelationships between IOs, agents, and auction outcomes by modeling IOs on a more granular level to retain their individual attributes as well as encoding their connections to other IOs and agents. This enables more expressive computations and a deeper understanding of the auction dynamics. Furthermore, we use a reward alignment technique [6] to optimize multiple KPI metrics simultaneously [7], forming a multi-objective optimization framework.
The contributions of this paper are twofold: first, we introduce a graph-based approach to auto-bidding that captures the complex, multi-agent interactions in auction environments; second, we propose the use of latent diffusion models for generating optimized bidding trajectories. Through extensive empirical evaluations on both synthetic [8] and real-world auction datasets [9], we demonstrate that our method significantly improves auto-bidding performance across several KPI metrics, as well as forecasting accuracy for auction outcomes.
Figure 2: Inverse Dynamics Model for Bid Computation.
II Latent Graph Diffusion Model for Auto-Bidding
@@
II-A Graph Embedding For Ad Auctions
The objective of the graph embedding is to represent the IOs presented to each agent in an expressive and comprehensive manner. In large-scale auctions, the number of IOs exposed to each agent can range drastically, from 00 to $10$s of thousands at every time-step. To process data with such variability yet interdependency, we represent these IOs as nodes of a graph and process their features using graph convolutions with an attention-based graph neural network (GNN) [10]. Rather than using a relational GNN, we instantiate two virtual nodes for each agent to connect exposed and non-exposed IOs with undirected links separately. By using bidirectional links between IOs and agent nodes, this design choice enables information propagation between agents with multiple hops. The additional links dedicated to non-exposed IOs aim to maintain context during sparse time periods. For scalability and computational limitations for larger scale simulations (e.g., dense graphs with $>1$ million edges), we perform random neighbor sampling on the non-exposed links to avoid prohibitively large computations, as opposed to a hierarchical graph building approach [11]. To compute the agent embedding, we aggregate the node embeddings of the two virtual nodes with a summation operation.
We optimize the latent graph embedding $x^{i}_{t}$ of agent $i$ at time-step $t$ generated by the GNN to contain predictive information regarding the bids $b_{t}^{i}[k]$ made for each IO $k$. Hence, we introduce a separate inverse dynamics model (IDM) to support the optimization of the following objective:
$\mathcal{L}_{\text{graph}}(\theta)=\mathbb{E}_{(t,k)}[\log\text{P}(b_{t}^{i}[k]|\theta,x_{t}^{i},x_{t+1}^{i},f_{t}^{i}[k],c_{t}^{i})]$ (1)
where $k$ is the IO identification index and $f_{t}^{i}[k]$ are the node embedding of the corresponding IO. To ensure the feasibility of this optimization, we append a context vector $c_{t}^{i}$ that provides more information regarding the agent agnostic to the individual IO, as shown in Figure 2. We optimize the objective over the parameters $\theta$ of both the GNN and the IDM. We view this optimization to (1) instill bid-policy information into the embedding and (2) constrain the graph embedding to reflect auction dynamics.
Figure 3: Agent $i$’s Approximation of Other Agents’ Graph $\omega^{i}(\mathcal{G}_{t}^{-i})$ by creating $h$ subgraphs used in student model $\mathcal{S}$.
@@ II-A Graph Embedding For Ad Auctions
The advertiser’s visibility of other agents’ IOs and bids is contingent on the auction design, e.g., sealed and closed envelope auctions, which could prevent a complete and accurate construction of the subgraphs of other agents. In such cases, we introduce a belief graph $\omega^{i}(\mathcal{G}_{t}^{-i})$, which approximates the other agents’ sub-graph by creating $h$ agent subgraphs, as shown in Figure 3. For our experiments, we set $h=4$. For simplicity, the nodes are randomly linked, where IO nodes are partitioned between the exposed and non-exposed nodes. To train the embedding, we apply knowledge distillation (KD) [12]. Therefore, we train a centralized GNN $\mathcal{T}$ using Equation 1 with full visibility of all agents’ sub-graphs. We then train a student model $\mathcal{S}$ using a response-based KD loss to mimic the outputs of $\mathcal{T}$.
@@
As defined so far, the embedding vector does not account for the temporal nature of the auction dynamics. Using a recurrent model representation or stacking multiple time steps are common approaches to handle such discrepancies. However, such methods are not scalable, nor can they handle open environments (e.g., advertisers can enter and leave the auction) trivially. Hence, we extend representation learning techniques, namely MA-SPL [13], to embed multi-agent and temporal dynamics information within the latent graph embedding $\{x^{i}_{0},x_{1}^{i},\dots,x_{T}^{i}\}_{i\in A}$, where $A$ is the set of all agents.
Figure 4: Computing Joint Embedding Vector using Self-Attention Network.
Equilibrium Computation
To promote a more cohesive multi-agent bidding strategy, we compute a joint embedding space by merging all agents’ graph embeddings. We refer to this design choice as equilibrium computation (EC), as the computation of all agents (i.e., bid computation and planning) is rooted in a shared latent state. Hence, we introduce an aggregation operation $\mathcal{A}:\{x^{i}_{t}\}_{i\in A}\mapsto x_{t}$ using a self-attention network, as shown in Figure 4. The self-attention network follows the Transformer encoder with a global mean pooling. This latent embedding $x_{t}$ is now shared between all agents when computing the generative objectives and for individual bid formulation.
II-B Latent Diffusion Models
The latent diffusion model (LDM) processes a temporal sequence of graph embeddings $\{x_{t}\}_{t\in[\tau,\tau+T]}$, where $T$ is the length of the context window, and forms the foundation to devise a planning-based auto-bidding solution. In this work, we adopt the Decision Diffuser framework [14], denoising state trajectories and offloading the action generation to the IDM. Unlike recent works using transformer-based architectures [15] and score-based prediction [16], we found greater empirical success using a more traditional 1D temporal convolution network and denoising with $\epsilon$-prediction. As shown in Figure 1, time-invariant features of the agent, such as the agent’s category and total budget, are processed as conditional inputs. We optimize the diffusion objective using in-painting masks [17].
@@ II-B Latent Diffusion Models
To ensure that the LDM’s posterior aligns with the KPI metrics and auto-bidding objective, we follow the procedure set in [6], utilizing an iterative sequence of constrained alignment optimization of reinforcement learning, direct preference optimization, and supervised fine-tuning with rejection sampling. We train a value function using IQL [18] to predict not only the value, i.e., the number of total conversations, but also other important KPI metrics, such as CPA, ROI, win rate, budget adherence, and social welfare [8].
  • Cost per acquisition (CPA) - CPA is the ratio between the total cost incurred and the earned value earned over all IOs, measuring the general performance of the auto-bidding strategy.
  • Return on Investment (ROI) - ROI measures the profitability of the strategy, computed as the ratio between the net profit and the total cost.
  • Win Rate - Win rate measures the effectiveness of a bidding strategy or decision-making process in securing wins and is computed as the ratio between the number of winning bids and the number of total bids made.
  • Budget Adherence - Budget adherence measures how well the spending aligns with the allocated budget. This metric ensures that resources are used efficiently and within the expected limits and is computed as the exceedance rate of the budget spending.
  • Social Welfare - Social welfare represents the aggregate value or utility generated by the system, measured as the total earned value over all advertisers.
With this value function, we can also strictly enforce KPI constraints, e.g., budget adherence, through rejection sampling.
@@
Training LGD-AB occurred in two stages. We first trained the graph embedding module and then the LDM separately. We then train both modules together, fine-tuning the GNN at a lowered learning rate. We found that having separate training stages was much more stable, and the additional stage of joint training benefits the performance of both modules. For partial observable settings, we train the teacher model using the described approach and then train the student model afterward.
III Results and Ablations Studies
Table I: Comparing performances between DiffBid and LGD-AB ($\mathcal{S}$, EC, and post-alignment) with mean and ±$\pm 1$ std over $64$ random seeds.
@@ III Results and Ablations Studies
III-A Experiment Setup
We evaluate LGD-AB on two auto-bidding simulations: AuctionNet [9] and synthetic auction [8].
AuctionNet
AuctionNet [9] contains over $500$ million in real-world records of online advertising instances, with $21$ separate bidding periods of $48$ time-steps and $48$ different advertisers. We use an auto-bidding simulator, following prior works [19], which reenacts auction dynamics using the dataset by having the auto-bidding agent impersonate one of the $48$ advertisers.
Synthetic Auction
Synthetic auction, on the other hand, generates artificial auction data that mimics real ad auctions by capturing their underlying properties through hierarchical feature generation and specific probabilistic distributions. We extend a prior work by Deng et al. [8] with configurable parameters of IOs to introduce longer and variable ad life cycles, multi-slot winning IOs, as well as a budget constraint on each agent. We collect an offline RL dataset of $100$ thousand auction simulations with a varying number of advertisers, allocation rule (e.g., FPA, GSP, VCG), bidding period durations, and IO’s properties mentioned above. Within this dataset, advertisers randomly select a uniform bid-scaling strategy or a non-uniform bid-scaling algorithm with pre-computed parameters [8].
III-B Empirical Evaluation
Table II: Auction Forecasting Results of LGD-AB Design Choices.
$\mathcal{L}_{fc}(\theta)$
AuctionNet Synthetic Auction
$\mathcal{T}$ $3.43$ $2.59$
$\mathcal{S}$ $2.64$ $2.02$
No EC $3.37$ $2.13$
EC $3.32$ $2.62$
No Alignment $3.25$ $2.09$
Alignment $3.02$ $2.04$
Figure 5: KPI Alignment Learning Curve on Synthetic Auction with multi-KPI alignment (shown in red) compared to optimizing each KPI criterion independently (shown in blue). For social welfare, we compare the training between EC and non-EC variants, as shown in the dotted lines.
Auction Forecasting
$\mathcal{L}_{fc}(\theta)=\mathbb{E}_{x,t}[\log\text{P}(x_{t^{\prime}>t}|\theta,x_{t^{\prime}\leq t})]$ (2)
In Table II, our evaluation was performed using our online simulators, where to compute $F(\theta)$, we use a testing dataset that was not used during training. We found that even with KD only, the student model retains, on average, $77.48\%$ of the teacher model’s forecasting accuracy despite the limitations placed by the partial observability. EC also improved forecasting capability in Synthetic Auction; however, in AuctionNet, it failed to replicate this improvement. We attribute this failure to the limited number of joint data samples (i.e., only $21$ bidding periods in total). To assist with this limitation and mitigate the efforts from data scarcity, we curated and trained on synthetic data generated by the LDM, utilizing a quality-based rejection sampling. By training on this synthetic dataset, we found a more fair representation of its forecasting accuracy, achieving a score of $0.21$ to $3.32$. Lastly, the KPI alignment reduced average forecasting accuracy by $9.47\%$, which is expected, as such fine-tuning may lower the likelihood of lower-value auto-bidding outcomes.
@@ III-B Empirical Evaluation
We compare our LGD-AB framework to another diffusion-based auto-bidding solution, DiffBid [5], across six common KPI metrics. We highlight that the main differences of the DiffBid framework to our LGD-AB are its use of a heuristic feature space, bid-scaling, and a conditional return input. In the original work of DiffBid, the authors have made comparisons to other competitive generative models (i.e., Decision Transformer) and RL baselines (i.e., BC, CQL, and IQL) to DiffBid, where DiffBid has demonstrated to outperform these other methods on similar auto-bidding simulations.
From Table I, we observe improvements across all KPIs using our LGD-AB framework from DiffBid. The results demonstrate a significant increase in the number of IOs won, shown by an average ×$\times 1.29$ improvement in return from our competitive baseline. This was achieved while significantly maximizing the win rate, lowering CPA, and maintaining budget adherence. Additionally, our proposed LGD-AB has demonstrated better social welfare, indicating an overall higher collective utility to the auction. In both our baseline and LGD-AB solution, budget adherence was not an issue and remained comparable, maintaining near-perfect budget adherence over all trials. We note that budget adherence was applied as a constraint via rejection sampling with a learned function with fair empirical risk and was not optimized using the alignment methods.
In Figure 5, we find that optimizing over multiple KPI (i.e., return, CPA, ROI, win rate, and social welfare) at the same time provides meaningful improvements within the individual performances of each metric. For some metrics, the difference is more significant, such as return and ROI, which are quite important measures in real-world applications.
Lastly, while EC did not show significant improvements over all KPI metrics, we found its utility in enabling much healthier alignment of social welfare.
Table III: Comparing Bid Computation Accuracy of DiffBid and LGD-AB.
@@
Accuracy of Bid Computation
We compare the bid prediction accuracy of DiffBid and LGD-AB in Table III. The results show a noticeably lower $\ell_{2}$ error with our LGD-AB’s $\ell_{2}$, alluding to being able to extract bids more accurately from our graph embeddings than the pre-constructed feature heuristic and limitations of the bid-scaling auto-bidding approaches. Notably, in the Synthetic Auction simulation dataset, where portions of the auction data are generated from non-uniform bid-scaling strategies, we notice a wider margin of difference in accuracy in favor of our LGD-AB framework.
@@
We observe some key issues and limitations while developing our proposed solution, mainly pertaining to the scale of data and efficiency of our generative modeling algorithm. With AuctionNet, while there exists a large set of individualized auction records to train on, there was an insufficient number of collective auction records that span full bidding periods to properly train a generative model to capture the underlying auction dynamics accurately. We addressed this limitation with synthetic data generation and hyperparameter tuning on the diffusion model. However, this limited performance can be improved with advancements in either training generative models in low data regimes (e.g., few-shot learning) or greater data availability. Another issue concerns the computational limitation from scaling the graph representation w.r.t. the number of IOs and the number of advertisers at every time-step, which was handled using neighbor sampling, which limits expressivity and completeness in the embedding computation. In future works, a form of dynamic graph sparsification or hierarchical graph representations could mitigate this. Lastly, a key
IV Conclusion
In this work, we propose a latent diffusion planning framework using graph-based embedding for auto-bidding applications. We demonstrate the utility of a graph-based representation, which enabled the LDM to capture the underlying dynamics distribution of a multi-agent auction process. By aligning various KPIs with the trained LDM, we devised an effective bidding strategy that is capable of generating bids that maximize various KPI metrics.